5x/6-14=11
17-8x/9=-7
x/2-2/5=3/4

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

5x/6-14=11
Multiply thru by 6 to get:
5x - 84 = 44
5x = 128
x = 25 3/5
====================
17-8x/9=-7
Multiply thru by 9 to get:
153 - 8x = -63
8x = 216
x = 27
=================
x/2-2/5=3/4
Multiply thru by 20 to getP
10x - 8 = 15
10x = 23
x = 2.3
================
Cheers,
Stan H.
=================

Answer by mananth(16949)

(5x/6)-14=11
LCD = 6
multiply by 6
5x*6/6 -14*6 =11*6
5x-84=66
5x=150
x=30
...
all the problems are to be done in the same way.
2. multiply equation by 9
3. LCD of 2,4,5 =20. multiply equation by 20.
m.ananth@hotmail.ca